# EMG-Driven Stiffness-Modulating Palpation for Telerehabilitation

### v1

@@ III-A Results & Discussion
The processed EMG signals were mapped to desired pressure values ($p_{d}$), which were tracked by the PID-controlled vacuum system of HJ-Pal. As shown in Fig. 3, the measured internal pressure closely followed the EMG-derived pressure commands. The system achieved low root-mean-square error (RMSE) values of $1.006\textpm 0.031$kPa during supination and $0.904\textpm 0.056$kPa during pronation. Furthermore, the measured pressure correlated strongly with the EMG-derived commands (Pearson’s correlation with $r=0.994\textpm 0.0004$ in supination; $r=0.966\textpm 0.010$ in pronation, all $p<0.0001$), confirming that the HJ-Pal is capable of responding to the high-frequency, noisy nature of EMG input.

### v2

@@ III-A Results & Discussion
The processed EMG signals were mapped to desired pressure values ($p_{d}$), which were tracked by the PID-controlled vacuum system of HJ-Pal. As shown in Fig. 3, the measured internal pressure closely followed the EMG-derived pressure commands. The system achieved low root-mean-square error (RMSE) values of $1.006\textpm 0.031$kPa in supinator and $0.904\textpm 0.056$kPa in pronator. Furthermore, the measured pressure correlated strongly with the EMG-derived commands (Pearson’s correlation with $r=0.994\textpm 0.0004$ in supinator; $r=0.966\textpm 0.010$ in pronator, all $p<0.0001$), confirming that the HJ-Pal is capable of responding to the high-frequency, noisy nature of EMG input.